OpenStudy (anonymous):

What are the solutions for x in the proportion 2x+3/x = 2x/2

OpenStudy (anonymous):

Can you simplify this with common denominators ?\[0 = \frac{2x +3}{x} - \frac{2x}{2}\] than \[0 = \frac{(2)(2x+3)}{(2)(x)} - \frac{(x)(2x)}{(2)(x)}\]